On January 1, a company issues bonds dated January 1 with a par value if $490,000. The bonds mature in 5 years. The contract rate is 8%, and interest is paid seminannually on June 30 and December 31. The market rate is 9% and the bonds are sold for $470,600. The journal entry to record the second interest payment using the effective interest methond of amortization is:

A) Debit Interest Expense $21,247.96; credit Discount on Bonds Payable $1647.96; credit Cash $19,600.

B) Debut Interest Expense $18,023; debit Premium on Bonds Payable $1,577; credit Cash $19,600.

C) Debit Interest Expense $18,023; debit Discount on Bonds Payable $1,577; credit Cash $19,600.

D) Debit Interest Expense $21,177; credit Discount on Bonds Payable $1,577; credit Cash $19,600.

E) Debit Interest Payable $19,600; credit Cash $19,600
Answer #1

Par value of bonds = $490,000

Sale value = $470,600

Discount on bonds payable (490,000 - 470,600) = $19,400

Contract rate = 8%

Market rate = 9%

First interest payment.

Interest payment = $490,000 * 8%*6/12 = $19,600

Interest expense = $470,600 * 9%*6/12 = $21,177

Discount on bonds payable = $1,577

So, Bonds payable at end of first interest payment = $470,600 + $1,577 = $472,177

Second interest payment.

Interest payment = $490,000 * 8%*6/12 = $19,600

Interest expense = $472,177 * 9%*6/12 = $21,247.96

Discount on bonds payable = $1,647.96

So, Journal entry to record second year interest payment is

Interest expense account...........Debit $21,247.96

To Discount on bonds payable...Credit $1,647.96

To Cash account........................Credit $19,600

Option 'A' is correct

Debit interest expense $21,247.96, Credit discount on bonds payable $1,647.96, credit Cash $19,600